A solution is made by dissolving 8 grams of a salt in 100 cm³ of water (density = 1 g/cm³). Assuming the volume does not change, what is the density of the solution?

A) 1.00 g/cm³

B) 0.92 g/cm³

C) 1.08 g/cm³

D) 1.04 g/cm³

E) 0.08 g/cm³


1
Expert's answer
2021-02-15T02:46:46-0500

Total mass of solution = 100 + 8 = 108 g

Volume of solution = 100 cm3

Density = 108/100 => [1.08] cm3